Suppose a consumer has income of 200 birr per month and he wants to spend all of his income on two goods, X and Y, whose prices are birr 5 respectively. Based on this information answer the following questions;
A. Express the budget line of the consumer both algebraically and diagrammatically.
B. Compute the equation of the budget line and interpret the result
C. Determine the slope of the budget line and interpret the result
D. What will happen to the original budget line?
I/ if money income doubles
II/If the price levels increases by 50%
III/ If Price of good Y doubles
1
Expert's answer
2020-01-27T10:10:01-0500

`Budget line expressed algebraically:

M=pxQx+pyQy

Diagrammatically



The equation of the budget line

M=PxQx+PyQy

Qy=M/py-Px/Py .Qx

M=income

Py=price of y

Px=price of x

Qy=200/5-5/5x

Qy=40-1x

Qy=40-x

The slope of the budget line

slope=-(Px/Py)

0-40/40-0

-40/40

=-1

What will happen to the original budget line

  • if money income doubles the original budget line will not move .Exactly the same quantities can be purchased as before
  • if the price levels will increase then the budget line will move inwards
  • if the price of y doubles the line would bivot outward on the point where the budget line crosses the horizontal axis